Can you help me with two:how do you change 2x+5y=15 and x-2y=9 into slope intercept form
Answer by algebrahouse.com(1659)

Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept

2x + 5y = 15
5y = -2x + 15 {subtracted 2x from both sides}
y = (-2/5)x + 3 {divided both sides by 5}

x - 2y = 9
-2y = -x + 9 {subtracted x from both sides}
y = (1/2)x - 9/2 {divided both sides by -2}

For more help from me, visit: www.algebrahouse.com